OpenStudy (anonymous):

Sara invested $8,000, part at 6% and the rest at 9% interest per year. How much did she invest at each rate if her total return per year on the investment was $660?

OpenStudy (anonymous):

Set up a system of equations using the information you have. If you let X = the amount invested at 6% per year, and Y= the amount invested at 9% per year, then X+Y = 8000. You can then set up the equation: X*.06+Y*.09=660 From these, you should be able to solve for X and Y

OpenStudy (anonymous):

I still don't get it.. @Vandreigan

OpenStudy (anonymous):

You can solve one equation for one of the variables and sub it into the other equation, finding an actual value for the other variable. For example, from X + Y = 8000 we get that X = 8000 - Y. Sub that in for X in the second equation and solve for Y.

OpenStudy (anonymous):

Thank you :) @Vandreigan

OpenStudy (anonymous):

My pleasure